# Latent heat estimation with machine learning

### v2

@@ Abstract
Tables 1 and 2 present estimates of critical energies $e_{o}$ and $e_{d}$ and latent heat for the 10-component Potts model, PM-10. Each row of critical energy estimates is followed by a row with the ratio of the deviation of the estimated critical energy to the statistical error to show the quality of the estimate. We find no noticeable finite-size corrections.
@@
In the case of a continuous transition driven by thermodynamic fluctuations, we see that finite-size analysis of machine learning probabilities reflects these fluctuations and leads to a reasonable estimate of the critical length exponent [3, 4]. In the case of a discontinuous phase transition, the correlation length does not diverges, and the dependence on the finite size may be different. It is known that finite-size corrections to thermodynamic quantities in the Potts model are very sensitive to the way they are estimated and depend on the quantities of interest. In two-dimensional space, the corrections can be proportional to $1/L$ or $1/L^{2}$ or even $1/L^{4}$ [25, 26, 24, 19].
@@
The phase prediction probability of a neural network is not a thermodynamic function, but at the same time its fluctuations can somehow reflect thermodynamic fluctuations, which may lead to the visible finite-size corrections. The correlation length is finite and approximately equal to 10.6 for the PM-10 model [24]. Interestingly, even moderate sizes of the studied systems compared to the correlation length allow us to estimate the critical values of the energies, and through them the latent heat, with an accuracy not worse than few percent.
@@
Table 1: Estimates of the critical energies and latent heat for PM-10 using raw data set RS.
@@
$\cal L$ 0.682(22) 0.677(30) 0.704(5) 0.697(10) 0.696049…
$\Delta\cal L$0.014 -0.019 0.008 0.001
Table 2: Estimates of the critical energies and latent heat for PM-10 using majority/minority data set MS.
Tables 3 and 4 present estimates of critical energies $e_{o}$ and $e_{d}$ and latent heat for the 20-component Potts model, PM-20. The quality of estimates is as good as those in the case of the PM-10 model. Again, our data show no noticeable dependence of the estimates on the system size. In this case, the correlation length in the critical point is much smaller at about 2.7 [24], and yet we see no finite-size corrections in our estimates. In contrast, in our recent study on PM-10 and PM-20 [19] using the Wang-Landau algorithm in simulations and the same MCPA algorithm as in the present research, the finite-size corrections to the critical energies estimated from the energy probability distribution show $1/L$ corrections, as predicted by the analytics [25, 26].
@@
Table 3: Estimates of the critical energies and latent heat for PM-20 using raw data set RS.
@@
$\cal L$ 1.186(3) 1.186(29) 1.164(38) 1.183(11) 1.19415…
$\Delta\cal L$0.008 -0.008 -0.030 -0.011
Table 4: Estimates of the critical energies and latent heat for PM-20 majority/minority data set MS.
Figure 5: Enlarged details of the left panel of the figure 4 showing small probability values with some features discussed in the Discussion section. Potts model with 20 components, PM-20 with grid size $L=60$.
Discussion. — In contrast to binary classification in the case of a phase transition of the second kind, our approach is based on ternary classification, using for training exactly known energies of the ordered phase $e_{o}$ and disordered phase $e_{d}$. To do this, we need to train the neural network on samples modeled at a certain value of energy $e$. For this purpose, we use the microcanonical population annealing algorithm (MCPA) [16, 17, 19], which anneals a large population of the modeled system in energy space. The trained network is used to classify all samples taken at a given energy $e$ into those that are more likely to be in ordered, disordered, or mixed phase. Thus, we can estimate the probability that a given sample belongs to one of the three phases. This protocol allows us to estimate critical energies and latent heat with reasonable accuracy.
At the same time, we found that the estimated probabilities contain some information about the details of the mixed phase. It is widely believed that there are four phase transitions in the mixed phase, which is a random mixture of ordered and disordered phases [14, 16]. In the case of the Potts model, these are droplets that reflect the ordered and disordered phases [27].
In the figure 5, the solid orange line corresponds to the mixed phase probability $P_{MS}$, the dotted green line corresponds to the disordered phase probability $P_{DS}$, and the dashed blue line corresponds to the ordered phase probability $P_{OS}$. The rightmost peak of $P_{OS}$ can be associated with the transition at energy $e_{1}(L)$ using the notations of Rose and Machta [16]. It is associated with the fluctuating droplets of OS phase within the DS sea, which vanishes at $e_{d}$. It is the precursor of the phase transition from mixed phase MS to disordered phase DS, as $e_{1}(L)\rightarrow e_{d}$ in the thermodynamic limit. By analogy, the leftmost peak of $P_{OD}$ can be associated with the transition at energy $e_{4}(L)$ and reflects DS droplets within the emerging sea of OS phase. Again, $e_{4}(L)\rightarrow e_{o}$ and the DS droplets completely vanishes at $e_{o}$. So, the transitions at $e_{1}(L)$ and $e_{4}(L)$ can be obtained only in the systems of finite size, and are not real phase transitions.
More interesting is the presence of small extrema on the DS and OS curves in the middle of the mixed phase. They can be related to the wrapping clusters when DS or OS droplets reach opposite boundaries of the system and due to the periodic boundaries will “wrap” around the torus. It is argued [27, 14, 16] that the transitions at $e_{2}$ and $e_{3}<e_{2}$ associated with the wrapping droplet OS and wrapping droplet DS, respectively, exist in the thermodynamic limit. Note that the extrema shown in our figure 5 are qualitatively similar to the extrema of a very different function, the integrated autocorrelation time shown in the Fig. 5 of the article [16].
@@ Acknowledgements.
The simulation was carried out using the high-performance computing resources of the National Research University Higher School of Economics. The research was supported by the Russian Science Foundation Grant No. 22-11-00259.

### v3

@@ IV.2 Energy and latent heat estimations
Tables 2 and 3 present estimates of critical energies $e_{o}$ and $e_{d}$ and latent heat for the 10-component Potts model, PM-10. Each row of critical energy estimates is followed by a row with the ratio of the deviation of the estimated critical energy to the statistical error to show the quality of the estimate. The estimates are scattered around a precisely known value of the critical energy. We found no noticeable finite-size dependence.
@@
In the case of a continuous transition driven by thermodynamic fluctuations, we see that finite-size analysis of machine learning probabilities reflects these fluctuations and leads to a reasonable estimate of the critical length exponent [3, 4]. In the case of a discontinuous phase transition, the correlation length does not diverges, and the dependence on the finite size may be different. It is known that finite-size corrections to thermodynamic quantities in the Potts model are very sensitive to the way they are estimated and depend on the quantities of interest. In two-dimensional space, the corrections can be proportional to $1/L$ or $1/L^{2}$ or even $1/L^{4}$ [34, 33, 26, 19]. A possible explanation for the apparent lack of finite size dependence of probability is that, for some reason, the amplitude $A$ of the correction to the probability function, which, by analogy with the energy probability density function, should behave as $A/L$ [33], is small. Nevertheless, we will see below that thermal fluctuations do depend on the lattice size, as can be seen from the behaviour of the variation $D_{xS}(E)$ of predictions.
@@
The phase prediction probability of a neural network is not a thermodynamic function, but at the same time its fluctuations can somehow reflect thermodynamic fluctuations, which may lead to the visible finite-size corrections. The correlation length is finite and approximately equal to 10.6 for the PM-10 model [26], see also Table 1. Interestingly, even moderate sizes of the studied systems compared to the correlation length allow us to estimate the critical values of the energies, and through them the latent heat, with an accuracy not worse than few percent.
@@
Table 2: Estimates of the critical energies and latent heat for PM-10 using raw data set RD.
@@
[…]
$\Delta\cal$ L -0.014 -0.019 0.008 0.001
Table 3: Estimates of the critical energies and latent heat for PM-10 using majority/minority data set MD.
Tables 4 and 5 present estimates of critical energies $e_{o}$ and $e_{d}$ and latent heat for the 20-component Potts model, PM-20. The quality of estimates is as good as those in the case of the PM-10 model. Again, our data show no noticeable dependence of the estimates on the system size. In this case, the correlation length in the critical point is much smaller at about 2.7 [26], and yet we see no finite-size corrections in our estimates. In contrast, in our recent study on PM-10 and PM-20 [19] using the Wang-Landau algorithm in simulations and the same MCPA algorithm as in the present research, the finite-size corrections to the critical energies estimated from the energy probability distribution show $1/L$ corrections, as predicted by the analytics [34, 33].
@@
Table 4: Estimates of the critical energies and latent heat for PM-20 using raw data set RD.
@@
$\Delta\cal$ L -0.008 -0.008 -0.030 -0.011
Table 5: Estimates of the critical energies and latent heat for PM-20 majority/minority data set MD.
IV.3 Finite-size rounding of the variation of the phase predictions
In Imry’s paper, it was shown [28] that the rounding off of some thermodynamic functions in the case of the first-order phase transition is inversely proportional to the square of the system size times the latent heat. Lee and Kosterlitz [33] (see also [34]) calculated this rounding in the more detail for the case of the Potts model, and, for example, the rounding of the specific heat is proportional to the logarithm of the number of states, divided by the latent heat and
$\Delta T\approx\frac{T_{c}\ln{q}}{{\cal L}L^{2}},$ (4)
and amaizingly, the difference with Imry’s formula is only in the multiplier with the number of components, which of course was outside Imry’s analysis, which considered the general case of the first-order phase transition due to the thermal fluctuations and Landau theory [27].
Analysis of the rounding in other quantities, such as Binder cumulants of energy and magnetization in the Potts model [33], shows that the terms in the expression (4) remain for all thermodynamic quantities, and the expression can be modified by additional multipliers such as the ratio of critical energies. Thus, it seems that the expression (4) is generally universal for the observables in the Potts model.
Figure 6: Variations $V_{xS}(E)$ of the probabilities of phases $P_{xS}(E)$ for $L=30$ (first row), 40 (second row), 50 (third row) and 60 (last row) for 10-state Potts model, PM-10. Left panel is the training/testing with the raw dataset RD and right panel is the training/testing with the majority/minority dataset MD. Blue dashed lines are ordered variations $P_{OS}$, orange solid lines are coexistence variations $P_{CS}$, and green dashed lines are disordered variations $P_{DS}$. The vertical lines denote the exact values of the critical energies of the ordered and disordered phase.
Figure 7: Same as in Fig. 6 for 20-state Potts model, PM-20.
As a result of the machine learning we estimates probabilities $P_{xS}$ of the phases (3). Using parallel with the specific heat which is proportional to the variation of the energy, we can analyze variation $V_{xS}$ of the phase probabilities $P_{xS}$ (remind the reader that $xS$ stands for one of the three phases - OS, CS, and DS)
$V_{xS}=\frac{1}{N_{test}}\sum_{i=1}^{N_{test}}\left[p_{xS}^{i}(e)\right]^{2}-\left[\frac{1}{N_{test}}\sum_{i=1}^{N_{test}}p_{xS}^{i}(e)\right]^{2}.$ (5)
Figures 6 and 7 show the variation $V_{OS}$, $V_{CS}$, and $V_{DS}$ of the probabilities of the three phases as functions of energy for different lattice sizes for PM-10 and PM-20. The variation maximum occurs around the critical energies $e_{d}$ and $e_{0}$, with fluctuations around $e_{d}$ are more pronounced. We use a Gaussian approximation in the neighborhood of the maximum, thus estimating the position of the maximum $\mu$ and its width $\sigma$. The position of the maximum $\mu$ gives estimates of the critical energies given in Tables 6-9, which are compatible with the estimates from the probabilities and given in Tables 2-5.
[…]
First, $\sigma$ depends on the linear lattice size $L$ inversely proportional to $1/L^{2}$. This is illustrated in Figures 8 and 9, where $\sigma$ is plotted against $1/L^{2}$ and the solid lines correspond to the fits to data from the last columns in Tables 6-9.
Second, we can calculate the ratio of the PM-10 peak width to the PM-20 peak width, namely $\sigma_{10}(L)$$\sigma_{20}(L)$. To do this, we compute the ratio of the width $\sigma_{10}(L)$ for a particular lattice size $L$ from the last column of Table 6 to the width $\sigma_{10}(L)$ for a particular lattice size $L$ from the last column of Table 8 for the MD dataset. Similarly, we compute the ratio for the RD dataset. The resulting ratio is shown in the Table 10. The ratio of the specific thermal widths for PM-10 and PM-20 calculated by the (4) method is approximately 1.318. All ratios are of the order of one, from which we can conclude that the variation width of the $V_{CS}$ probability does depend on the number of spin components and the latent heat. What we have not been able to determine is the influence of other numbers characterizing a particular Potts model with a particular number of spin components. We mentioned in this subsection that a particular quantity can have different multipliers. There are many of them, as mentioned in [33].
In any case, we verify that the probability width $V_{CS}$ does not contradict the general form (4).
Dataset L $\mu$ $e_{o}$ $\sigma$
MD 30 -1.66374(12) -1.664252 0.01308(14)
40 -1.66852(5) 0.00920(6)
50 -1.66635(6) 0.00727(8)
60 -1.66614(3) 0.00606(4)
RD 30 -1.65281(17) -1.664252 0.02937(21)
40 -1.66555(10) 0.01805(11)
50 -1.66668(6) 0.01493(7)
60 -1.66233(4) 0.01217(5)
Table 6: Estimation of peak position $\mu$ and width $\sigma$ from the variation $V_{CS}$ near energy $e_{o}$, for PM-10 model.
Dataset L $\mu$ $e_{d}$ $\sigma$
MD 30 -0.99447(97) -0.968203 0.07163(138)
40 -0.99953(81) 0.05723(87)
50 -0.97043(32) 0.05268(60)
60 -0.97458(23) 0.04555(37)
RD 30 -0.95612(47) -0.968203 0.04773(67)
40 -0.97693(21) 0.02918(27)
50 -0.95333(16) 0.02437(21)
60 -0.97422(10) 0.01893(12)
Table 7: Estimation of peak position $\mu$ and width $\sigma$ from the variation $V_{CS}$ near energy $e_{d}$, for PM-10 model.
Dataset L $\mu$ $e_{o}$ $\sigma$
MD 30 -1.81636(8) -1.820684 0.00926(9)
40 -1.82305(7) 0.00636(9)
50 -1.82014(3) 0.00474(4)
60 -1.81827(3) 0.00393(4)
RD 30 -1.81825(16) -1.820684 0.02153(20)
40 -1.81859(9) 0.01606(11)
50 -1.81233(6) 0.01290(7)
60 -1.82047(4) 0.01003(5)
Table 8: Estimation of peak position $\mu$ and width $\sigma$ from the variation $V_{CS}$ near energy $e_{o}$, for PM-20 model.
Dataset L $\mu$ $e_{d}$ $\sigma$
MD 30 -0.66018(146) -0.626529 0.08682(202)
40 -0.65383(96) 0.06924(130)
50 -0.66662(96) 0.05614(90)
60 -0.64900(58) 0.05680(75)
RD 30 -0.61982(66) -0.626529 0.06752(118)
40 -0.65190(57) 0.04730(75)
50 -0.60936(24) 0.03381(33)
60 -0.63806(12) 0.02692(16)
Table 9: Estimation of peak position $\mu$ and width $\sigma$ from the variation $V_{CS}$ near energy $e_{d}$, for PM-20 model.
Figure 8: Width $\sigma_{10}(L)$ of the variation $V_{CS}$ with error bars from Tables 6 and 7 with the linear fit (solid line). Upper row represent data for the left peak in Figures 6 and bottom row represent data for the right peak in Figures 6. Left row is the RD data set and right row is the MD data set. Data sets of PM-10 model, $q=10$.
Figure 9: Width $\sigma_{20}(L)$ of the variation $V_{CS}$ with error bars from Tables 8 and 9 with the linear fit (solid line). Upper row represent data for the left peak in Figures 7 and bottom row represent data for the right peak in Figures 7. Left row is the RD data set and right row is the MD data set. Data sets of PM-20 model, $q=20$.
L left peak right peak
MD 30 1.41 0.825
40 1.45 0.827
50 1.53 0.938
60 1.54 0.802
RD 30 1.36 0.707
40 1.12 0.617
50 1.16 0.721
60 1.21 0.703
Table 10: Ratio of the peak width $\sigma_{10}(L)/\sigma_{20}(L)$ near the ordered energy $e_{o}$ and disordered energy $e_{d}$.
IV.4 Coexistence phase
In contrast to binary classification in the case of a phase transition of the second order, our approach is based on ternary classification, using for training exactly known energies of the ordered phase $e_{o}$ and disordered phase $e_{d}$. To do this, we need to train the neural network on samples modeled at a certain value of energy $e$. For this purpose, we use the microcanonical population annealing algorithm (MCPA) [16, 17, 19], which anneals a large population of the modeled system in energy space. The trained network is used to classify all samples taken at a given energy $e$ into those that are more likely to be in ordered, disordered, or coexistence phase. Thus, we can estimate the probability that a given sample belongs to one of the three phases. This protocol allows us to restore critical energies and latent heat with reasonable accuracy.
Figure 10: Enlarged details of the left panel, last line of the Figure 5 showing small probability values with some features discussed in the Discussion section. Potts model with 20 components, PM-20 with grid size $L=60$, RD dataset.
At the same time, we found that the estimated probabilities contain some information about the details of the coexistence phase. It is widely believed that there are four phase transitions in the coexistence phase, which is a random mixture of ordered and disordered phases [14, 16]. In the case of the Potts model, these are droplets that reflect the ordered and disordered phases [35].
In the Figure 10, the solid orange line corresponds to the coexistence phase probability $P_{CS}$, the dotted green line corresponds to the disordered phase probability $P_{DS}$, and the dashed blue line corresponds to the ordered phase probability $P_{OS}$. The rightmost peak of $P_{OS}$ can be associated with the transition at energy $e_{1}(L)$ using the notations of Rose and Machta [16]. It is associated with the fluctuating droplets of OS phase within the DS sea, which vanishes at $e_{d}$. It is the precursor of the phase transition from coexistence phase CS to disordered phase DS, as $e_{1}(L)\rightarrow e_{d}$ in the thermodynamic limit. By analogy, the leftmost peak of $P_{OD}$ can be associated with the transition at energy $e_{4}(L)$ and reflects DS droplets within the emerging sea of OS phase. Again, $e_{4}(L)\rightarrow e_{o}$ and the DS droplets completely vanishes at $e_{o}$. So, the transitions at $e_{1}(L)$ and $e_{4}(L)$ can be obtained only in the systems of finite size, and are not real phase transitions.
More interesting is the presence of small extrema on the DS and OS curves in the middle of the coexistence phase. They can be related to the wrapping clusters when DS or OS droplets reach opposite boundaries of the system and due to the periodic boundaries will “wrap” around the torus. It is argued [35, 14, 16] that the transitions at $e_{2}$ and $e_{3}<e_{2}$ associated with the wrapping droplet OS and wrapping droplet DS, respectively, exist in the thermodynamic limit. Note that the extrema shown in our Figure 10 are qualitatively similar to the extrema of a very different function, the integrated autocorrelation time shown in the Figure 5 of the article [16].
[…]
V Discussion
The application of machine learning in statistical physics is developing intensively. The main objective is to investigate the details of applicability of this new research method. In other areas of physics, such as high-energy physics [36], the use of a trained neural network model accelerates the pre-screening of uninteresting events in the analysis of particle scattering data. Training a neural network is a very labor-intensive process, but the trained network makes predictions at a much faster rate than other software models. Using this approach in statistical physics provides correct statistics when generating statistically correct system configurations, for example, using diffusion networks [37].
In our proposed ternary sample classification method, we get the opportunity to investigate the probability of belonging to one of the three phases of the system state. We are not aware of any other approach that allows us to estimate such probabilities. As we noted above,our approach allows us to examine the details of the coexistence phase.
It would be interesting to apply the same ternary classification approach to study phase transitions in non spin-system domain. For example, a fairly simple yet interesting problem with three phases is the behaviour of the hard disks in two-dimensions [38]. The solid, hexatic, and liquid phases are separated by a continuous transition and a first-order phase transition. It would be natural to use the density values $\eta_{sh}=0.700$ and $\eta_{hl}=0.716$ from the paper [38], which separates these phases, in the training process and use trained neural network model to predict the probabilities of all three phases.
@@ Acknowledgements.
The simulation was carried out using the high-performance computing resources of the National Research University Higher School of Economics. The first short version of this article was supported by Russian Science Foundation grant No. 22-11-00259, and the final extended version was supported by grant No. 25-11-00158.